OpenStudy (mattyboyy):

2. Cars in auto shops are raised up using hydraulic lifts that are controlled by large cylinders of different sizes that are connected by a chamber that holds fluid. Suppose a car sitting on a hydraulic lift exerts a downward force of 2,000 N on a piston with an area of 0.3m2. Calculate the pressure that the car exerts on the piston

OpenStudy (michele_laino):

hint: the pressure P exerted on that piston, is given by the subsequent computation: \[\Large P = \frac{F}{S} = \frac{{2000}}{{0.3}} = ...\frac{{Newtons}}{{{m^2}}}\]